OpenStudy (anonymous):

Find the solution(s): sqrtx+9+sqrtx=9

OpenStudy (anonymous):

I saw your name and had no choice but to answer this now...to bad I don't understand the question :s

OpenStudy (anonymous):

Heh

OpenStudy (solomonzelman):

\[\sqrt{x+9}+\sqrt{x}=9\]right?

OpenStudy (anonymous):

Yes

OpenStudy (solomonzelman):

square both sides, using \[(a+b)^2=a^2+2ab+b^2\]for the left side,\[(x+9)+(2 \times \sqrt{x+9} \times \sqrt{x})+x=81\]

OpenStudy (solomonzelman):

\[2x+9+2\sqrt{x^2+9x}=81\]

OpenStudy (solomonzelman):

\[2\sqrt{x^2+9x}=-2x-9+81\]\[2\sqrt{x^2+9x}=-2x+72\]\[2~~(\sqrt{x^2+9x})=2~~(-x+36)\]\[\sqrt{x^2+9x}=-x+36\]

OpenStudy (solomonzelman):

square both sides,\[x^2+9x=x^2-72x+1296\]\[81x=1296\]\[x=16\]

OpenStudy (anonymous):

Wow

OpenStudy (anonymous):

The reason why I'm confused with this is because my teacher taught it to me differently and I don't get how she got the answer. Can I show you the way she did it?

OpenStudy (solomonzelman):

sure.

OpenStudy (solomonzelman):

I'll be back in short...

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

The first thing she did was she subtracted sqrtx from the left side. Then she squared each side to get rid of the radical and I get x+9 = 81-x. Then I subtract 9 from both sides and add x to both sides and then divide both sides by 2 and get 36 but I already know that's the wrong answer.

OpenStudy (solomonzelman):

\[(9-\sqrt{x})^2~~~????????\]What a poor teacher, she forgot about the middle term.

OpenStudy (solomonzelman):

Even a 16 year old like me knows this, common teacher.....

OpenStudy (anonymous):

brb one sec